Let basket A be (x, y) = (4, 3). For each of the following utility functions (a-d), answer the following questions (iiv):

i) Determine the utility at basket A

ii) Write out the equation for the indifference curve that basket A lies on. (simplify to y equals some function of x)

iii) Find 3 other points that lie on the same indifference curve as A.

iv) Calculate the MRSx,y at basket A and give an economic interpretation of what the means in words.

(a) U (x, y) = 9xy.

(b) U (x, y) = 2x2 + 5y2 .

(c) U (x, y) = 6x1/2 + y.

(d) U (x, y) = 8x + 7y.
Answer #1

I have answered each parts except (iii) for functions (a-d)

The image below shows it.a) sve (i) A = (4,3) = 9x12 X 12Neらave MU (I ) utility = ax43 + 5x2 ごう子a. MO u t o adCL 2. iV 3 反 3ご 多2t 21 (01 ) C6 MRS

Sorry i forgot to add (ii) for (d)

equation for indifference curve is ,

8x + 7y = 53

7y = 53 - 8x

y = (53-8x) / 7

Hope this helps

please comment for doubts

thanks